Let: C = consumption, Ip = investment spending (as a function of price level), G = government spending, Tx = tax revenue, Yd = after-tax income, Assume for a given closed economy:

C=100 + 0.9 Yd – 20P

Ip= 400 – 40P

G=300

T=100

Moreover, aggregate supply curve for this economy is defined by the following equation:

P=1.41 + 0.0001Y

a.According to the investment equation (Ip= 400 – 40P) as overall price level in the economy increases investment spending decreases. How could you explain this situation? Please use graphs to elaborate your answer.

b. Find the equilibrium level of overall price and aggregate output in this economy. What would be the value of consumption and investment spending at this equilibrium?


1
Expert's answer
2020-06-29T14:44:35-0400

a) Higher prices, more expenditures, less saving, less investments


"Y=100+0.9*(Y-100)-20*(1.41+0.0001Y)+400-40*(1.41+0.0001Y)+300"

"Y=5900\ -\" the equilibrium level of output

"Ip=400-40*(1.41+0.0001*5900)=320"

"C=100+0.9*(5900-100)-20*2=5280"